$SPAD/src/input images2.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
-- Color gallery page 2
)r newton
)r cdraw
-- create a Newton's iteration function for the equation x**3 = 2.
f := newtonStep(x**3 - 2)
setClipValue(4)
drawComplexVectorField(f**3, -3..3, -3..3)
drawComplex(f**3, -3..3, -3..3)
drawComplex(f**4, -3..3, -3..3)

———-

References^{[1] nothing}

4